The length of a cylinder is measured with a metre rod having least count 0.1 cm. Its diameter is measured with vernier calipers having least count 0.01 cm. Given that length is 5.0 cm and radius is 2.00 cm. The percentage error in the calculated value of the volume will be:


        (a)    11%


        (b)    2%


        (c)    3%


        (d)    4%


Grade:10

1 Answers

Navjyot Kalra
10 years ago

(c)

    V = ?r2l

    ?V/V%=2 ?r/r%+?l/l%

    = 2 × 0.01/2×100%+0.1/5×100%=1%+2%=3%

?